OpenStudy (anonymous):

In ΔABC, if the lengths of sides a and c are 8 centimeters and 16 centimeters, respectively, and the measure of ∠C is 35°, what is the measure of ∠A to two decimal places?

OpenStudy (anonymous):

the choices are a. 16.67° b.17.50° c.26.57° d.35.53°

OpenStudy (anonymous):

never mind i couldnt spend anymore time on it so i had to guess

OpenStudy (anonymous):

@joejoe213 Assuming that a is the length of the side opposite the angle of measure A and that c is the length of the side opposite the angle of measure C, you can use the law of sines to get that sin(A)/a = sin(C)/c, so sin(A) = asin(C)/c = sin(35deg)/2. So A = arcsin(sin(35deg)/2) = 16.67.

OpenStudy (anonymous):

kk hi

OpenStudy (anonymous):

okay im about to leave school thanks anyway♥

OpenStudy (anonymous):

have a good thanksgiving yall!